OpenStudy (anonymous):

Given the equation y - 3 = one half(x + 6) in point-slope form, identify the equation of the same line in standard form. A.x - 2y = -12 B. y = one halfx C. y = one halfx + 6 D. y = one halfx + 9

OpenStudy (anonymous):

@Katelyn17

OpenStudy (anonymous):

Im guessing its c but im not sure

OpenStudy (anonymous):

was that it?

OpenStudy (anonymous):

the answer

OpenStudy (anonymous):

y - 3 = one half(x + 6) bring over the 3 and change the sign bro! y = one hallf(x + 6) + 3 multiplying out the stuff in brackets y = one half x + one half 6 + 3 y = onehalf x + 3 + 3 y = one half x + 6

OpenStudy (anonymous):

your answer is C bro!

OpenStudy (anonymous):

thx man your the best! :)